### v1

# Learning to accelerate distributed ADMM using graph neural networks

Abstract
Distributed optimization is fundamental in large-scale machine learning and control applications. Among existing methods, the Alternating Direction Method of Multipliers (ADMM) has gained popularity due to its strong convergence guarantees and suitability for decentralized computation. However, ADMM often suffers from slow convergence and sensitivity to hyperparameter choices. In this work, we show that distributed ADMM iterations can be naturally represented within the message-passing framework of graph neural networks (GNNs). Building on this connection, we propose to learn adaptive step sizes and communication weights by a graph neural network that predicts the hyperparameters based on the iterates. By unrolling ADMM for a fixed number of iterations, we train the network parameters end-to-end to minimize the final iterates error for a given problem class, while preserving the algorithm’s convergence properties. Numerical experiments demonstrate that our learned variant consistently improves convergence speed and solution quality compared to standard ADMM. The code is available at https://github.com/paulhausner/learning-distributed-admm.
1 Introduction
The emergence of large interconnected networked systems, where each agent in the network typically has access only to local private information, together with the recent explosion in data availability, has necessitated the development of efficient algorithms for distributed optimization across various domains such as machine learning, control, and signal processing. The alternating direction method of multipliers (ADMM) [22, 19] is well-suited for this task and has gained increased popularity due to its flexibility and scalability [8]. A key feature of ADMM is that it can be implemented in a distributed fashion, where the agents compute the iterations locally and in parallel. After computing their new iterates, the agents communicate the updates to their neighbors, and each node then updates its local variables by combining its own computations with the information received. While ADMM has a well-established convergence theory (see e.g. [16]), it is known to be highly dependent on the hyperparameters and may converge slowly to the global minimizer [42].
A promising avenue for addressing these limitations lies in recent advances in machine learning. In particular, the learning-to-optimize framework [12] aims to automate the design of faster algorithms by using data-driven techniques that enhance the convergence of iterative methods for a given problem class. Building on this idea, we adopt an unrolling scheme [30], where network parameters of our model are trained to minimize a loss function that reflects ADMM performance after a fixed number of iterations. In the context of distributed optimization problems, which consist of an interconnected network of agents, graph neural networks (GNNs) [7, 36] are a natural choice to parameterize the learning method as they can learn and reason about graph-structured data that can represent various problem domains such as interconnected network systems.
[…]
Figure 1: Computation structure of iteration $k$ of the decentralized distributed ADMM algorithm computed as two message-passing steps in our proposed GNN (see Algorithm 2.2). The initial node features ${V}^{k}$ contain the previous ADMM iterates $(x^{k}_{i},y^{k}_{i},\lambda^{k}_{i})$ for each node $i$ and the non-zero edge features E{E} describe the weighted pairwise connection of nodes in the network which is fixed across iterations. After two message-passing steps – each consisting of a message (msg), an aggregation (agg), and an update (upd) part indicated by a colored box in the figure – the new iterates $(x^{k+1}_{i},y^{k+1}_{i},\lambda^{k+1}_{i})$ of all nodes in the network are the output of the GNN.
In particular, in this paper, we highlight that classical decentralized and distributed ADMM iterations [29] can be implemented as a graph network using the recently popularized message-passing framework [7] as we showcase in Figure 1 and further expand upon in Section 4.1. Throughout this paper, we explicitly exploit this connection and utilize GNNs to accelerate the distributed ADMM algorithm. The contributions of our paper can be summarized as follows:
  • We show the one-to-one correspondence of distributed ADMM iterations [29] and message-passing networks [21], a popular framework to build graph neural networks. This shows a strong alignment [40] between ADMM and GNNs which we exploit in this paper.
  • Based on these insights, we develop and evaluate different approaches for learning to accelerate distributed ADMM through learning the hyperparameters of the algorithm corresponding to graph-, node-, and edge-level tasks to learn the step size and communication matrix.
  • Finally, we propose a network architecture and loss functions to train the graph neural network in an end-to-end fashion by unrolling the ADMM iterations.
In numerical experiments, we demonstrate that our learning-based approaches reduces the distance of the iterates to the minimizer compared to the baseline method and can improve the convergence of ADMM even beyond the unrolling steps used during training. A concise summary of the notation used in this paper can be found in Appendix A.
2 Related work
The goal of learning-to-optimize (L2O) approaches is to accelerate the convergence of parameterized optimization problems using machine learning techniques [1, 12]. In particular, we focus on accelerating the ADMM algorithm for distributed optimization problems; for a detailed overview of the ADMM algorithm we refer to the survey by Boyd et al. [8].
While L2O approaches can yield impressive results in practice [3], they typically lack convergence guarantees and may fail to generalize, especially on unseen problem instances or when the underlying data distribution shifts [1]. A popular approach to ensure convergence, which we also adopt, is to learn the hyperparameters of a classical optimization algorithm. Along those lines, Sambharya et al. [34] learn to warm start fixed-point algorithms, including ADMM, by unrolling the iterations and using automatic differentiation to compute the gradient through the iterations. In a follow-up work [33], they learn hyperparameters of fixed-point iterations. Similarly, Ichnowski et al. [28] use reinforcement learning to tune the hyperparameters of OSQP, an ADMM-based solver for convex quadratic programs. These approaches have also been combined to solve quadratic programming problems with distributed ADMM [35]. In contrast to this, we do not allow the presence of a central communication node. Further, we make no assumptions beyond the convexity of the objective function in our framework. A similar unrolling technique for ADMM iterations was successfully applied by Sjölund and Bånkestad [37] for nonnegative matrix factorization. Another approach by Noah and Shlezinger [31] learns the algorithms themselves directly through ADMM unrolling. We instead learn a network that predicts the step size hyperparameters of the optimization algorithm for a certain problem instance, either on a global or local level, increasing the flexibility of our learning method.
Further, we propose a novel way to also learn the communication matrix. The predicted weighted Laplacian matrix can be seen as a form of learned preconditioning that stabilizes the iterates by improving problem conditioning [38]. Machine learning methods previously have been successfully applied to learn preconditioners for linear systems [24, 25], imaging problems [18], and gradient descent [20].
Since the communication network and the ADMM iterations can be represented as a graph, graph neural networks (GNNs) [7] are a natural fit for parameterizing the unrolled iterations. GNNs have been put to other L2O uses previously, as they can represent various optimization problems, such as quadratic programs [13], and can simulate the iterations of existing solvers such as interior point methods [32]. However, while Qian et al. [32] show that GNNs can, in principle, approximate the solution mapping by simulating the solver iterations, no convergence guarantees are obtained.
Further, GNNs have also recently been applied to general distributed optimization. However, instead of starting with a learned algorithm and adding constraints inspired by theory [26], we formulate distributed ADMM as a GNN architecture operating on the communication network. Then, we parameterize the update functions with neural networks, allowing the GNN to learn on the given problem class, thereby replacing extensive manual hyperparameter tuning with a data-driven procedure.
3 Background
We start by briefly describing the distributed optimization problem setting and subsequently explain how ADMM iterations can be used to solve this problem. Finally, we give a short overview of the message-passing framework, which we later use to implement the ADMM iterations.
3.1 Distributed optimization
A distributed optimization problem consists of a network of $m$ interconnected agents, represented by the nodes $\mathcal{V}=\{1,2,\dots,m\}$ of a connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. An edge $(i,j)\in\mathcal{E}$ symbolizes the opportunity for agent $i$ and $j$ to communicate. The agents jointly aim to solve the optimization problem
$\min_{x\in\mathbb{R}^{n}}\quad\sum_{i=1}^{m}f_{i}(x)\,$, (1)
where the local objective function $f_{i}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ is only known to agent $i$ and its data parameterizing the objective cannot be communicated. Limited communication between agents can arise, for instance, due to limited networking capacity, as in edge networks [15], or due to privacy concerns [14]. In such scenarios, we must solve the problem in a distributed manner, decoupling the objective minimization and adding a consensus constraint:
$\min_{x_{1},x_{2},\dots x_{m}}\quad$ $\sum_{i=1}^{m}f_{i}(x_{i}),$ (2a)
s.t. $x_{1}=x_{2}=\dots=x_{m}\,$, (2b)
where each agent has its own local optimization variable $x_{i}\in\mathbb{R}^{n}$. Consensus among the agents can be reached by communicating these local solutions with their neighbors as defined by $\mathcal{G}$.
We assume that all $f_{i}$ are convex and that the objective (2a) is a proper, closed function. While there exist many different algorithms that solve this problem, we focus on distributed ADMM for peer-to-peer networks, i.e., the agents communicate directly with each other. In other words, we consider the decentralized setting, which is more general – and challenging – than having a central hub or intermediary.
3.2 ADMM
The alternating direction method of multipliers (ADMM) is an operator splitting technique that solves convex optimization problems of the form
$\min_{x,z}\;f(x)+g(z)\quad\text{s.t.}\quad Ax+Bz=c\,$, (3)
where $x\in\mathbb{R}^{p}$, $z\in\mathbb{R}^{q}$, $A\in\mathbb{R}^{r\times p}$, $B\in\mathbb{R}^{r\times q}$, and $c\in\mathbb{R}^{r}$. Originally proposed by Glowinski and Marroco [22] and Gabay and Mercier [19], ADMM combines features of dual ascent, allowing parallel updates, and the method of multipliers, which converges under mild assumptions [8].
Distributed ADMM
Following the approach proposed by Makhdoumi and Ozdaglar [29] we rewrite the distributed problem (2) to match the general ADMM formulation (3). In the rewritten form, the variable $x=[x_{i}]_{i=1}^{m}$ corresponds to the stacked individual variables of the agents and the function $f(x)=\sum_{i=1}^{m}f_{i}(x_{i})$ is the global objective. Further, the consensus constraint (2b) is enforced by the equivalent formulation $(P\otimes I_{n})x=0$, where $P$ is a $m\times m$ communication matrix satisfying $\text{null}(P)=\text{span}\{1_{m}\}$. This is in particular fulfilled by the (weighted) Laplacian $\mathcal{L}$ of a graph.
To allow parallel computation within ADMM, one can introduce an auxiliary variable $z$ and define $g$ as a convenient indicator function. Moreover, by exploiting the inherent symmetry of the optimization variables, the edge-based auxiliary variables $z_{ij}$ can be replaced with corresponding node-based variables $y_{i}$, while the original dual variables are reparameterized as related variables $\lambda_{i}$. This reparametrization reduces redundancy and leads to a more compact and efficient formulation [29].
ADMM iterations
In the resulting distributed ADMM algorithm one iteration consists of updating all variables $x_{i}$, $y_{i}$, and $\lambda_{i}$ for each of the agents $i$. The updates can be executed for each agent $i$ in parallel and require only information about the local objective and previous iterates of its neighbors in the graph denoted by $\mathcal{N}(i)=\{j:(i,j)\in\mathcal{E}\}\subset\mathcal{V}$.  [^1] We do not allow for self-loops, so $i\notin\mathcal{N}(i)$. Further, since $\mathcal{G}$ is undirected $(i,j)=(j,i)\in\mathcal{E}$. The local solution candidates $x_{i}$ contain the solution to a regularized version of the local subproblem of agent $i$ and are computed as the minimizer to
$x_{i}^{k+1}=\arg\min_{x_{i}}\left(f_{i}(x_{i})\,+\hskip-7.11317pt\underset{j\in\mathcal{N}(i)\cup\{i\}}{\sum}\hskip-7.11317pt\left((\lambda_{j}^{k})^{T}(P_{ji}x_{i})+\frac{\alpha}{2}\|P_{ji}(x_{i}-x_{i}^{k})+y_{j}^{k}\|_{2}^{2}\right)\right)\,$, (4)
where $\alpha>0$ is the global step size hyperparameter. After communicating the updated iterates computed in (4) to all neighbors, the ADMM updates for $y_{i}$ and $\lambda_{i}$ are given by:
$y^{k+1}_{i}=\frac{1}{d_{i}+1}\sum_{j\in\mathcal{N}(i)\cup\{i\}}P_{ij}x_{j}^{k+1},\qquad\qquad\lambda_{i}^{k+1}=\lambda_{i}^{k}+\alpha y_{i}^{k+1}.$ (5)
Note that the variable $y=[y_{i}]_{i=1}^{m}$ is a weighted residual for consensus since the constraint $(P\otimes I_{n})x=0$ implies consensus on the components of $x_{i}$. The full algorithm is given in Alg. 1.
Convergence
Under the assumptions on the objective function from section 3.1, the ergodic average of the computed local solutions $x_{i}$ converges for every global step size $\alpha>0$ and any communication matrix $P$ with sublinear rate $\mathcal{O}(1/k)$ against the global solution. However, both the global step size $\alpha$ and the communication matrix $P$, affect the convergence speed in practice, and optimal choices appear to be instance-dependent [29]. The convergence results can be extended to the case where there exists not one but a sequence of step sizes [16], or even node-level step sizes $\alpha_{i}>0$ [6]. Moreover, the solution to (4) does not need to be exact to ensure convergence [16].
3.3 Graph neural networks
Graph neural networks (GNNs) learn a latent feature representation for each node in a graph based on some input features and the underlying graph topology. In this paper, we focus on the framework of message-passing neural networks (MPNNs) which update the node features of a graph by aggregating information from the neighboring nodes and edges [21]. Given edge weights $e_{ij}$ for $(i,j)\in\mathcal{E}$ represented by the weighted adjacency E{E} of the graph and initial node features $v_{i}^{0}$ for each node $i\in\mathcal{V}$, the latent node features $v_{i}^{l+1}$ are iteratively updated in each message-passing step $l$ by computing
$v^{l+1}_{i}=\texttt{update}\left(v^{l}_{i},\texttt{aggregate}(\{\texttt{message}(e_{ij}\cdot v^{l}_{j}:j\in\mathcal{N}(i))\})\right).$ (6)
Here, the update and message functions can contain parameters that are learned during training, which are denoted by $\theta$. We denote the node feature matrix in layer $l$ with ${V}^{l}$. The aggregate function is typically permutation invariant, as there is no ordering of the graph neighbors, and can be implemented, for example, using the sum or mean function. This makes the overall architecture permutation equivariant. Thus, the ordering of the nodes does not affect the output of the network.
This framework can be easily extended to more general update and message functions, including global features and allowing nodes to send messages to themselves through self-loops [7].
4 Method
The key observation of our paper is that the distributed ADMM iterations introduced in Section 3.2 can be rewritten in the form of two message-passing steps of form (6). In the following, we explicitly establish the connection between the ADMM iterations and GNNs in Section 4.1. Further, we describe how the hyperparameters of the iterations can be parameterized using neural networks. In Section 4.2, we then describe how to train the parameters of this network using unrolled ADMM iterations.
4.1 Learning ADMM via message-passing networks
We start by deriving a one-to-one mapping from each distributed ADMM update step in (4) and (5) to a corresponding message-passing step. Each step consists of a respective message, aggregate, and update function. This correspondence is visualized in Figure 1.
Input representation
The distributed optimization problem can be directly encoded as the input to a GNN. The underlying graph structure is naturally given by the communication network $\mathcal{G}$ from the problem formulation (2). The node features of the graph consist of the three ADMM iterates $(x_{i},y_{i},\lambda_{i})$. Furthermore, the non-negative edge weights $e_{ij}>0$ of the graph represent the weighted connectivity, that determine the strength of the connection between two nodes in the graph. To obtain a valid communication matrix from the edge features, we construct the weighted Laplacian $P=\tilde{D}-E$ where $\tilde{D}$ is a diagonal $m\times m$ matrix of the weighted degree for each node $\tilde{D}_{ii}=\sum_{j\in\mathcal{N}(i)}e_{ji}$ and $E$ the symmetric matrix representation of the graph connectivity.
Message-passing steps
We now explain our reformulation of an ADMM iteration in form of two message-passing steps. Recall that in the MPNN framework in (6), the node update function only has access to the aggregation of the incoming messages and not the individual messages, as required in the original ADMM update step (4). To overcome this difficulty and define a valid update function, we rewrite (4) equivalently to
$x_{i}^{k+1}=\underset{x_{i}}{\text{arg min}}\,f_{i}(x_{i})+\bigl{(}P_{ii}\lambda_{i}^{k}+\bar{\lambda}_{\rightarrow i}^{\,k}+\alpha(P_{ii}y_{i}^{k}+\bar{y}_{\rightarrow i}^{\,k})\bigl{)}^{T}x_{i}+\frac{\alpha}{2}\hskip-5.69046pt\sum_{j\in\mathcal{N}(i)\cup\{i\}}\hskip-14.22636pt\|P_{ji}(x_{i}-x_{i}^{k})\|_{2}^{2}.$ (7)
In the rewritten form, $(\bar{\lambda}_{\rightarrow i}^{k},\bar{y}_{\rightarrow i}^{k})$ is the summation (agg) of all to node $i$ incoming messages that we define by $(P_{ji}\lambda_{j}^{k},P_{ji}y_{j}^{k})$. Hence, in accordance with the MPNN framework, the refined update function in equation (7) now only depends on the aggregated features and the nodes’ local information that is encoded in the node features. For details of the derivation, we refer to Appendix D.
The updates for $y$ and $\lambda$ in equation (5) can be carried out jointly within an second message-passing step, and their translation is straightforward. It suffices to define the message between two adjacent nodes as $P_{ij}x_{j}^{k+1}$ and replace the respective part in (5) with the aggregate of these messages.
Together, this yields a two-step message-passing network that we visualize in Figure 1 and summarize in Appendix D in Algorithm 2.2. We highlight the step for the $x$-update by a red box, and the $y$- and $\lambda$-update by a dashed blue box. Due to the equivalence to the original iteration in (4) and (5), the network’s output coincides with the updated iterates of a single distributed ADMM iteration.
Learned components
The message-passing steps derived in the previous paragraph rigidly reproduce the decentralized distributed ADMM iterations, without any added flexibility through machine learning. In other words, there are no learnable parameters that can be utilized to improve the performance of the iterations. In the following, we present several ways to introduce learnable parameters in the message-passing steps that still maintain convergence under assumptions stated in Section 3.1. These approaches correspond to different levels of tasks for the GNN:
  • Graph-level task: We learn a global step size by letting every node predict an $\alpha_{i}$ based on the local information and using the averaged $\alpha=\frac{1}{m}\sum_{i=1}^{m}\alpha_{i}$ for each node. In the distributed setting, this corresponds to introducing additional communication to reach consensus over the step size.
  • Node-level task: Instead of choosing the same step size $\alpha$ for each node, we can learn an individual local $\alpha_{i}$ for each node. This avoids additional communication between the nodes since the averaging step is omitted. However, each node is required to choose $\alpha_{i}$ without knowledge about the step sizes of the other nodes, making it conceptually more challenging.
  • Edge-level task: For each edge in the network, we can learn the positive edge weight $e_{ij}$ between two connected agents. Based on the edge weights predicted before the first ADMM iteration, we use the corresponding weighted Laplacian $P$ as a fixed communication matrix across all steps.
For the first two tasks, the step size is predicted using a small neural network that receives as input the optimization variables of the previous iteration, consisting of the primal variable $x$, and the variables $y$ and $\lambda$ for each node. Further, we append the aggregated messages and the total number of nodes in the graph to the input. We apply an instance normalization layer before passing the input through the network to ensure numerical stability [39]. At the beginning of the algorithm, we can predict and change the step size in every iteration, resulting in a separate MLP for every change. However, to ensure convergence, we fix the predicted step size after a predefined number of $L$ iterations. [^2] In practice, we often choose $L=K$.
For the predicted edge weights, we use the local degree profile [10], describing the local connectivity of the graph, as an input to an MLP. The feature representations of the two nodes get stacked and passed through the learned network to predict a positive edge weight that preconditions the ADMM iterations. To ensure invariance to the node order and symmetrize the edge weight graph, we sum the predicted edge weights for each directed edge. To ensure the positivity of the learned algorithm hyperparameters $e_{ij}$ and $\alpha$, we use the softplus activation function on the network output.
4.2 Network training
In real-world applications, distributed optimization problems need to be solved repeatedly for similar communication networks $\mathcal{G}$ and similar local objectives $f_{i}$, providing potential for problem-class specific acceleration. We assume access to a dataset $\mathcal{D}$ consisting of a sequence of problem instances that share an underlying structure to train our model to improve the average performance across the dataset.
Unrolling
We apply a technique called algorithm unrolling [12], where the algorithm is trained on a fixed number of iterations to maximize performance within this iteration budget. In our case, we construct a GNN, consisting of $K$ distributed ADMM iterations via the message-passing scheme defined in the previous section. Thus, in total we use $2K$ message passing steps for the GNN implementation.
Depending on the chosen learnable components, several MLPs are incorporated into the GNN to influence the update functions. The parameters of all the neural networks are denoted by $\theta$. The final output contains the approximate solution to the optimization problem (2) which now depends on these parameters. The model is then trained using standard gradient-based learning methods and learns how to choose the learnable components (by adjusting $\theta$) so that a suitable loss function is minimized.
Optimization subproblem
There is one obvious difficulty in making the GNN fully differentiable. In each update of the node features $x_{i}$, an optimization problem of the form (7) needs to be solved, and the gradient of the network parameters $\theta$ needs to be computed with respect to the solution to this optimization problem. In the case of step size learning, all of $\alpha_{i}$, the previous iterates, and incoming messages depend on the parameters of the network through previous updates. To still compute the gradient, we need to differentiate the solution to the optimization problem with respect to the parameters of the network $\frac{\partial x_{i}^{k+1}}{\partial\theta}$ which can be decomposed via the chain rule $\frac{\partial x_{i}^{k+1}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial\theta}$.
Computing the first part can be achieved in closed form by the means of implicit differentiation [2]. However, implicit differentiation can in practice be costly and requires additional overhead. Instead, we solve the subproblem (7) approximately by unrolling an iterative solver [^3] Here, we unroll the iterative solver within each unrolled iteration of ADMM. or use the closed-form solution to (7) if it exists. In both cases, we compute its gradient using standard automatic differentiation techniques. Along the same lines, the partial derivative of the optimization problem (4) with respect to the learned edge weights $e_{ij}$ is computed. The remaining update steps can be differentiated using standard automatic differentiation tools provided by JAX [9].
Loss function
The loss function is designed such that it evaluates the GNN outcome $x^{K}=[x_{i}^{K}]_{i=1}^{m}$ only after the $K$ unrolled iterations of ADMM. Thus, we do not care if the learned ADMM iterations perform worse than the baseline method after $k<K$ steps. For practical problems, however, the performance for more iterations of ADMM is also evaluated during testing.
To train our model, we utilize the dataset $\mathcal{D}$. Each problem instance from $\mathcal{D}$ consists of a graph $\mathcal{G}$ and the parameterization of the per-node local objective functions $f_{i}$. Additionally, we have access to the true minimizer $x^{\star}_{d}$ that solves problem (2) for instance $d\in\mathcal{D}$ and can be computed offline efficiently as the problem is convex. Further, we precompute $\hat{x}^{K}_{d,i}]_{i=1}^{m}$, the approximate solution obtained by running the baseline distributed ADMM iteration with the default step size and communication matrix. This is used to normalize the loss function in (8) and avoid over-weighting more involved instances in the training data. We train the network to minimize the empirical risk of the training data by minimizing the loss function
$\ell(\theta;\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{d=1}^{|\mathcal{D}|}{}\frac{1}{m_{d}}\left(\sum_{i=1}^{m_{d}}\frac{\|x_{d,i}^{K}(\theta)-x^{\star}_{d}\|_{2}^{2}}{\max\left(\|\hat{x}_{d,i}^{K}-x_{d}^{\star}\|_{2}^{2},\varepsilon\right)}\,\right),$ (8)
using stochastic gradient-based optimization and automatic differentiation. Here, $m_{d}$ is the number of agents in the problem instance $d\in\mathcal{D}$, and $\varepsilon$ is a small positive constant that prevents division by zero.
This loss function is regression-based as it relies on the true solution to the optimization problem (2) that can be computed apriori and is part of $\mathcal{D}$ [1]. Further, the loss function only cares about minimizing the distance to $x^{\star}_{d}$ without directly training the network to fulfill the consensus constraint (2b) or to minimize the global objective (2a). The denominator does not depend on the network parameters but is used to normalize the loss functions across problem instances. Moreover, the normalized loss allows a natural interpretation of the network performance as it measures the network outcome relative to the baseline method with fixed hyperparameters.
5 Experiments
5.1 Problem setting
We present and evaluate our method on two different use-cases of the distributed ADMM algorithm which are inspired by real-world problems. More details about the problem generation for the problems can be found in Appendix B.
Network average consensus problem
The first problem we consider is network consensus where all agents have some local information $b_{i}$ and the goal for the agents is to find the mean of all information across the network [43]. Despite its simplicity, this problem is a classic benchmark that highlights some of the important design choices in distributed optimization. The objective function for each agent $i$ in the network can be encoded as:
$f_{i}(x_{i})=\|x_{i}-b_{i}\|^{2}.$ (9)
Distributed least-squares problem
For the second use-case we consider an application area inspired by machine learning where we want to solve a distributed least squares problem [41]. Here, each node has access to only a part of the data given by inputs $B_{i}$ and corresponding labels $b_{i}$ and the goal of the agents is to collaboratively find the least-squares solution xx. [^4] In the machine learning literature, the data is usually referred to as $(x,y)$ pairs and the parameters of the linear regression model as $\theta$. In order to be consistent with the ADMM framework presented above, we use $x$ as the regression parameters and $(B_{i},b_{i})$ as the data. Thus, the local objective is:
$f_{i}(x_{i})=\|B_{i}x_{i}-b_{i}\|^{2}.$ (10)
5.2 Results
We train our model for $K=10$ unrolling steps to minimize the loss function (8). Additional details about the model training can be found in Appendix C. In Table 1, we show the results of different learned methods in terms of their ability to minimize the loss function (8) and achieve consensus among the iterates across the network.
In the middle column, we show the results obtained for $K=10$ iterations which is the value used for the loss function in (8). We can see that all learned methods outperform the baseline, achieving a lower error across all learned methods in both experimental settings. This also holds for consensus among the iterates, even though the loss only influences this indirectly. Overall, the combined method clearly performs best in both metrics.
Table 1: Comparison of different methods of the ADMM algorithm after a fixed number of K steps. Numbers reported are the average across the test data. The error measures the average distance of the local solutions to the global minimizer $\texttt{avg}\left(\|x_{i}^{K}-x^{\star}\|^{2}_{2}\right)$ and the consensus measures the degree to which the constraint (2b) is satisfied as $\texttt{avg}\left(\|x_{i}^{K}-\bar{x}^{K}\|\right)$, where $\bar{x}^{K}$ is the average solution at step $K$. The baseline method does not use learning but uses the default values $\alpha=1,e_{ij}=1$ for all iterations instead. In the combined version, we learn both individual step-size parameters $\alpha_{i}$ and the edge weights simultaneously. Lower is better for all metrics. Best is bold, second base underlined.
k=5 K=10 (train) k=20
Consensus Error Consensus Error Consensus Error Consensus
Baseline 19.07 12.23 9.17 7.10 2.33 2.13
Global $\alpha$ 18.74 17.21 4.19 3.65 3.46 3.37
Local $\alpha_{i}$ 14.50 12.26 3.05 2.82 1.17 1.14
Weight $e_{ij}$ 17.98 10.96 7.39 5.47 1.52 1.40
Combined 13.39 8.05 1.96 1.76 0.76 0.65
Least-squares Error Consensus Error Consensus Error Consensus
Baseline 70.86 10.29 53.23 7.77 30.09 2.44
Global $\alpha$ 61.32 23.41 27.12 9.21 14.93 3.48
Local $\alpha_{i}$ 56.80 14.66 23.79 7.37 12.96 1.84
Weight $e_{ij}$ 67.83 14.19 43.99 8.20 19.45 2.51
Combined 58.97 8.83 18.24 5.42 11.79 0.95
Additional iterations
In the leftmost column of Table 1, we compare the results for $k=5$ steps. We can see that generally, the gap between the baseline method and the learned methods is not as big as for the training objective. However, already after $5$ iterations, the learned methods outperform the baseline in most cases, even though the training only evaluates the performance after $K=10$ steps.
In the rightmost column, we show the performance of the (learned) ADMM algorithm after $k=20$ iterations. For the additional iterations, we use a constant step size $\alpha=1$ as in the baseline. This allows us to guarantee the convergence of the ADMM iterations by fixing the step size for all remaining iterations. Moreover, we apply the same learned weighted Laplacian matrix in the additional iterations. Also in this situation, apart from the global step size in the consensus problem, we observe advances compared to the baseline. Unsurprisingly, the combined method shows the strongest improvements.
In summary, although our method is trained to perform well only after the fixed $K=10$ iterations, both earlier and later iterates still outperform the baseline method in our experiments.
Function value
In addition to the comparisons in the previous paragraph, we also evaluate the obtained results in terms of the objective function value. We measure this by the relative objective at iteration $k$, which is the difference in the objective values normalized by the minimal value:
$\frac{\left|f(x^{k})-f(x^{\star})\right|}{|f(x^{\star})|}\,.$ (11)
In Figure 2, the relative objective is plotted for $k=20$ iterations for the different methods. We can clearly see that the performance of the learned methods is best after the $K=10$ unrolling steps for training. The model using a global $\alpha$ even shows a large initial increase in the objective when additional iterations are executed. Further, we can see that most methods initially perform worse than the baseline method but performance improves with more iterations.
One exception is the learned communication matrix for the least-squares problem which leads to a higher objective value for most of the iterations even though the distance to the minimizer gets reduced as we have seen in Table 1. However, when combining the communication matrix with the learned step size, the performance significantly improves also in terms the objective value showing that the preconditioning effect still benefits the ADMM convergence.
6 Discussion & conclusion
In this paper we have seen how the distributed ADMM algorithm can be accelerated by expressing the updates in form of message-passing steps. In the experimental section we utilize this to learn the step size and communication matrix but there are several other ways to further exploit this connection. In the following we highlight some of these potential extensions and discuss limitations of our approach.
Limitations
With the training scheme outlined in Section 4.2, we train the unrolled algorithm for a fixed number of iterations in our experiments to achieve a low loss. However, ADMM is also known to suffer from slow convergence when getting closer to the minima. With the current training scheme we only train the network to perform well after the $K=10$ unrolled steps which does not necessarily translate into a faster global convergence [4]. Further, we implement the differentiation of the optimization problem using unrolling of the inner solvers which has a larger memory footprint than implicit differentiation and as such limits the number of unrolled ADMM iterations we can train the network on. The specific method is a trade-off between computational time and memory requirements that can be chosen based on the underlying problem class and its dimensionality [35]. However, in practical real-time applications a certain computational budget typically exists, justifying our approach of tuning performance to a fixed number of steps.
Future work
In our approach, we only augment distributed ADMM by learning components that predict the hyperparameters. However, it is also conceivable to replace some of the steps completely. For example, the update of the $x$-variable requires solving a – possibly time-consuming – optimization problem. While it is tempting to replace this with the output of a neural network, it is significantly more challenging to ensure convergence for such inexact methods. Nevertheless, it has been successfully demonstrated for other optimization schemes [5]. Further, the exchange of messages involving all iterates can be computationally expensive. Accordingly, the framework can be extended to only communicate compressed information about the iterates without losing convergence [11].
Further, in our paper we focus on the decentralized distributed ADMM version presented by Makhdoumi and Ozdaglar [29] while many other formulations of distributed ADMM exist that can leverage GNNs. In particular GNNs with a global node can be directly mapped to distributed ADMM versions with a centralized communication hub that is connected to all nodes in the graph opening up for many more synergies between distributed optimization and graph neural networks in the future.
Figure 2: Semi-log plots of the relative objective for the two problem classes across all 2020 iterations. The network is only unrolled for the first $10$ iterations during training as in the previous results.
Conclusion
In this paper, we demonstrate an equivalence between distributed ADMM and message-passing networks that, to our knowledge, has not been explicitly recognized before. Based on this insight, we describe a learning-to-optimize approach that predicts instance-specific hyperparameters for distributed ADMM, which is trained by unrolling the algorithm and minimizing the normalized distance between the true minimizer and the network output for a specific problem distribution. These learning tasks can be naturally expressed as different graph learning tasks leveraging the connection between ADMM and GNNs. In numerical experiments, we validate that the learned algorithm outperforms the baseline method and leads to an accelerated convergence of ADMM.
Acknowledgments
The authors thank Daniel Arnström for helpful discussions and suggestions. This research was funded in part by Sweden’s Innovation Agency (Vinnova), grant number 2022-03023, the Centre for Interdisciplinary Mathematics at Uppsala University (CIM), and supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Glossary of Notation
Table 2: Summary of the notation used throughout the paper.
Symbol Description Comments
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ Graph with set of vertices $\mathcal{V}$ and edges $\mathcal{E}$ $\mathcal{V}=\{1,2,\dots,m\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$
$\mathcal{N}(i)$ Set of neighboring nodes of $i$ By default we assume $i\notin\mathcal{N}(i)$
$d_{i}$ Degree of node $i$ Equal to $|\mathcal{N}(i)|$
$D,\tilde{D}$ Degree matrix and weighted degree matrix
$m$ Number of agents / nodes in the graph
$n$ Dimension of the local optimization problem
$f_{i}$ Local objective function of node / agent $i$
$b_{i},B_{i}$ Local objective function information
$x_{i}$ Local optimization variable of node / agent $i$
$I_{n}$ Identity matrix of size $n$
$1_{n}$ Vector of all ones of size $n$
${V}=[v_{i}]_{i=1}^{m}$ Node features in GNN
${E}=[e_{ij}]$ for $(i,j)\in\mathcal{E}$ Edge features in GNN Typically implemented as a sparse matrix
$l$ Layer in GNN Twice as many GNN layers as ADMM steps
$\theta$ All learnable network parameters
$\mathcal{D}$ Dataset used for training
$\varepsilon$ Small constant to avoid division by zero
$\alpha$ ADMM step size
$\mathcal{L}$ Weighted Laplacian matrix $\mathcal{L}=D-E$
$P$ Communication matrix
$x,z$ ADMM primal variables Stacked and slack distributed variables
$y$ ADMM distributed variable This is a reformulation of the $z$ variable
$\lambda$ ADMM dual variable
$A,B$ ADMM linear constraint matrices
$k$ Iterations of the ADMM algorithm
$K$ Total number of ADMM steps in the unrolling
$L$ Number of steps to switch to fixed step size Often $L=K$
Appendix B Problem details
Here, we describe in more detail how we generate the different problem classes used in the experiments in Section 5 and describe how to generate ground-truth labels for them which are used for the supervised training as well as provide details on how the subproblem in equation (4) is solved in practice. Further, we provide the additional information to describe how the distribution of problems is generated for each problem.
For each problem class we generate a total of $900$ instances for training, $100$ instances for validation and $100$ instances for testing purposes. All results in the main text are evaluated on the testing instances which are not seen by the network during training.
B.1 Network average consensus
For the network average consensus problem, each node or agent $i\in\{1,\ldots,m\}$ in the network has a local objective function as in (9) to minimize the distance between the local optimization variable $x_{i}$ and its local private data $b_{i}$.
Problem generation
This local objective function is parameterized for every agent by a vector $b_{i}\in\mathbb{R}^{n}$. It is sampled independently from a $n$-dimensional normal distribution $\mathcal{N}(0,\Sigma)$ with covariance matrix $\Sigma=100I_{n}$. In our experiments, we fix $m=8$ and $n=2$.
Subproblem solution
The solution to the subproblem (4) for this use-case is a special case of the result presented for the distributed least-squares problem below with $B_{i}=I_{n}$. Since the resulting linear system for the problem that we derive in equation (13) is diagonal, the problem can be solved more efficiently in practice requiring only $\mathcal{O}(n)$ steps and does not require expensive matrix inversion techniques. Instead, we can take the reciprocal of the diagonal elements to obtain the inverse. Note that these elements are always non-zero by the construction of the linear system, and thus the inverse is well-defined.
Global solution
For training, we assume access to the global solution $x^{\star}$ for all problem instances. We can efficiently generate these solutions by utilizing the underlying problem structure and the direct access to all problem data. The minimizer is given by
$x^{\star}=\frac{1}{n}\sum_{i=1}^{n}b_{i}$ (12)
In other words, the agents are collaborating to find the mean of their vectors $b_{i}$ without sharing the actual data.
B.2 Distributed least squares problem
In the distributed least squares problem, the local objective function is of the form (10). In addition to the data vector $b_{i}$ from the network average consensus problem, each agent also has a private feature transformation matrix $B_{i}\in\mathbb{R}^{n\times n}$. The latter is used to weight the linear regressionc efficients of $x_{i}$ before comparing with $b_{i}$.
Problem generation
The local objective function is parametrized for every agent by the pair $(B_{i},b_{i})\in\mathbb{R}^{n\times n}\times\mathbb{R}^{n}$. Each entry of the matrix $B_{i}$ is sampled from a continuous uniform distribution $\mathcal{U}[0,1]$ on the unit interval. In addition, we assume in our experiments that $B_{i}$ is of full rank to obtain better convergence behavior of ADMM and ensure closed-form solutions. We enforce this by discarding and resampling $B_{i}$ in case the magnitude of an eigenvalue is below $1/10$. The data vector $b_{i}$ is sampled as before.
Across all problem domains, the underlying graph structure is based on the Erdős-Rényi random graph model [17] with $m=8$ nodes. We discard sampled graphs which are not connected.
Subproblem solution
In the distributed least squares problem, the solution to the optimization problem to update the $x$-variable from (4) has a closed form solution and can be solved as the solution to the following linear equation system:
$x_{i}^{k+1}=(2B_{i}^{T}B_{i}+\alpha M_{ii})^{-1}\bigl{(}2B_{i}^{T}b_{i}-P_{ii}\lambda_{i}^{k}-\bar{\lambda}_{\rightarrow i}^{k}+\alpha(M_{ii}x_{i}^{k}-P_{ii}y_{i}^{k}-\bar{y}^{k}_{\rightarrow i})\bigl{)},$ (13)
where $M_{ii}$ is a diagonal matrix defined by
$M_{ii}:$ $=\left(\sum_{j\in\mathcal{N}(i)}P_{ji}^{T}P_{ji}+P_{ii}^{T}P_{ii}\right)\otimes I_{n}$ (14)
$=\Biggl{(}\sum_{j\in\mathcal{N}(i)}e_{ji}^{2}+\bigl{(}\underbrace{\sum_{j\in\mathcal{N}(i)}e_{ji}}_{\text{weighted degree of $i$}}\bigl{)}^{2}\Biggl{)}\otimes\,I_{n}$
since $P_{ii}=\tilde{D}_{ii}=\sum_{j\in\mathcal{N}(i)}e_{ji}$ is the weighted degree in the case where we choose the weighted Laplacian as our communication matrix for the algorithm. Note that the (weighted) degree and the sum of the weighted connections can be easily computed as a message in the MPNN.
We can observe in particular that the matrix here is positive definite since $B_{i}^{T}B_{i}$ is positive (semi)-definite and $M_{ii}$ is a diagonal matrix with positive diagonal entries. Therefore, we can use the iterative conjugate gradient method to compute the solution to the equation system which scales computationally better than inverting or factorizing the matrix for large-scale problems.
Global solution
The global minimizer is given by
$x^{\star}=\left(\sum_{i=1}^{m}B_{i}^{T}B_{i}\right)^{-1}\left(\sum_{i=1}^{m}B_{i}^{T}b_{i}\right)\,.$ (15)
Note that the inverse exists since again $B_{i}^{T}B_{i}$ is positive definite for every $i$, due to the construction of the training data, and hence is the sum described in the problem generation paragraph.
Appendix C Training details
In the following, we discuss the details of the training scheme and the chosen hyperparameters in order to ensure reproducibility of our presented results.
Implementation details
We use the JAX framework [9] to implement the differentiable distributed ADMM algorithm as a message-passing network. In particular, we are using the jraph toolkit to implement the message-passing steps [23] and integrate it with the flax library to design neural network components using the novel linen API [27].
Network architectures
For the network components to predict the hyperparameters, within the message-passing steps, we are using a 2-layer multi-perceptrons. The hidden layer has $32$ neurons and there is $1$ output which gets passed through a softplus activation function to ensure positivity. We use a ReLU non-linearity between the two layers.
  • We use a single network for all iterations from $k=2$ until $K=10$. For the first step, we do not use learning but use the default step size. Thus, there are a total of $9$ learned steps in the unrolling scheme. The input to the network is given by the stacked optimization variables of the current iteration. Thus, the total number of trainable parameters is 3 $3\thinspace 75$3.
  • The input for the edge weight prediction is given by the concatenated local degree profiles of the adjacent nodes [10]. Each of these consists of $5$ features containing statistics about the the neighborhood structure including the degree, maximum and minimum neighboring degree, as well as the mean and variance. To ensure that the communication matrix is symmetric, we sum up the edge predictionsWe use the same predicted communication matrix for all iterations, thus, the number of parameters in the model is $385$.
  • In the combined model, we use both the step size network and the edge weight network together. Thus, the total number of parameters is $4\thinspace 138$.
Loss function
To avoid division by zero, we include the hyperparamter $\varepsilon$ into the loss function in (8). In the experiments, we set $\varepsilon=10^{-5}$.
Training hyperparameters
For the training, we are using the Adam optimizer with a learning rate of $10^{-4}$ and global gradient clipping of with radius $1.0$. We are training the model for $100$ epochs using a batch size of $5$. Thus, during training a total of $18\thinspace 000$ parameter update steps are executed.
Computational resources
All reported experiments are executed on a single NVIDIA TITAN Xp GPU with 12 GB of memory. Training a single epoch requires approximately $10$ minutes plus additional time for the validation process. In total, the training of one model for $100$ epochs requires around $20$ hours.
Appendix D Algorithms
D.1 Decentralized, distributed ADMM
In the following, we state the decentralized, distributed ADMM algorithm (Algorithm 1) from Makhdoumi and Ozdaglar [29] adapted to our notation. As described in section 3.2, each iteration updates all three iterates $(x_{i},y_{i},\lambda_{i})$ and includes two communication steps to pass information about the current iterates to the neighbors in $\mathcal{G}$.
Algorithm 1 Decentralized, distributed ADMM
1:Hyperparameters: $\alpha>0$, communication matrix $P$ ($\mathcal{L}$ here), $d_{i}$ degree of node $i$
2:Initialize $x_{i}^{0}\in\mathbb{R}^{n},\lambda_{i}^{0}=0\in\mathbb{R}^{n}$ for $i=1,\ldots,m$, and $k=0$
3:Communicate $x_{i}^{0}$ to all neighbors $j\in\mathcal{N}(i)$
4:for $i=1,\ldots,m$ do
5:  Initialize $y_{i}^{0}$
$y_{i}^{0}=\frac{1}{d_{i}+1}\sum_{j\in\mathcal{N}(i)\cup\{i\}}P_{ij}x_{j}^{0}$
6:end for
7:while stopping criterion is not reached do
8:  for $i=1,\ldots,m$ in parallel do
9:   Communicate $y_{i}^{k}$ and $\lambda_{i}^{k}$ to all neighbors $j\in\mathcal{N}(i)$
10:   Update local solution candidate $x$:
$x_{i}^{k+1}=\underset{x_{i}}{\text{arg min}}\,\,\left(f_{i}(x_{i})+\sum_{j\in\mathcal{N}(i)\cup\{i\}}\left((\lambda_{j}^{k})^{T}(P_{ji}x_{i})+\frac{\alpha}{2}\|P_{ji}(x_{i}-x_{i}^{k})+y_{j}^{k}\|_{2}^{2}\right)\right)\,.$
11:   Communicate $x_{i}^{k+1}$ to all neighbors $j\in\mathcal{N}(i)$
12:   Update local deviation-variable $y$:
$y_{i}^{k+1}=\frac{1}{d_{i}+1}\sum_{j\in\mathcal{N}(i)\cup\{i\}}P_{ij}x_{j}^{k+1}\,.$
13:   Update local dual-variable $\lambda$:
$\lambda_{i}^{k+1}=\lambda_{i}^{k}+\alpha y_{i}^{k+1}\,.$
14:  end for
15:  $k\leftarrow k+1$
16:end while
17:Return: $x^{k}=[\,x_{i}^{k}\,]_{i=1}^{m}$
D.2 Message-passing formulation of distributed ADMM
When it comes to defining a message-passing network that performs one iteration of Algorithm 1, it is critical to reformulate the update steps in such a way that they only require knowledge of the aggregation of incoming messages and not of all individual messages. We pointed this out in section 4.1 and defined respective message, aggregate, and update functions that respect the MPNN framework. The two resulting message-passing steps are presented in Algorithm 2.2. To highlight the one-to-one correspondence, we compare it with the equivalent Algorithm 2.1, which is identical to a single iteration of the original distributed ADMM algorithm (see lines of 8-14 of Algorithm 1).
Equivalence between updates.
For completeness, we will briefly show the equivalence between the original $x$-update step from (4) and the rewritten step in (7) here. Details for the $y$ and $\lambda$ update are omitted since the equivalence follows immediately from the definition of the messages.
Equation (4) (or line 8 of Algorithm 1) includes a sum over features from neighboring nodes that we need to revise. The first summand in this sum can be rewritten easily by the definition of the aggregated messages $\bar{\lambda}_{\rightarrow i}^{\,k}$:
$\sum_{j\in\mathcal{N}(i)\cup\{i\}}(\lambda_{j}^{k})^{T}(P_{ji}x_{i})=(P_{ii}\lambda_{i}^{k})^{T}x_{i}+\Biggl{(}\underbrace{\sum_{j\in\mathcal{N}(i)}P_{ji}^{T}\lambda_{j}^{k}}_{=:\,\bar{\lambda}_{\rightarrow i}^{\,k}}\Biggl{)}^{T}x_{i}\,$, (16)
Note that $P_{ji}$ must be included in the message since weighting is no longer possible once all $\lambda_{j}^{k}$ are aggregated. The second summand, the penalty term for consensus, splits into the following
$\sum_{j\in\mathcal{N}(i)\cup\{i\}}\|P_{ji}(x_{i}-x_{i}^{k})+y_{j}^{k}\|_{2}^{2}=\sum_{j\in\mathcal{N}(i)\cup\{i\}}\Biggl{(}\|$ $P_{ji}(x_{i}-x_{i}^{k})\|_{2}^{2}$
$\underbrace{+\|y_{j}^{k}\|_{2}^{2}+2(P_{ji}(x_{i}-x_{i}^{k}))^{T}y_{j}^{k}}_{\text{includes features from neighbors}}\,\Biggl{)}\,.$ (17)
Since the $x$-update optimizes the equation with respect to $x_{i}$, we understand that the quadratic term $\|y_{j}^{k}\|_{2}^{2}$ is irrelevant for this update. The mixed term, which involves features from neighboring nodes, splits as follows
$2\sum_{j\in\mathcal{N}(i)\cup\{i\}}(P_{ji}(x_{i}-x_{i}^{k}))^{T}y_{j}^{k}=2(x_{i}-x_{i}^{k})^{T}\Biggl{(}P_{ii}^{T}y_{i}^{k}+\underbrace{\sum_{j\in\mathcal{N}(i)}P_{ji}^{T}y_{j}^{k}}_{=:\,\bar{y}_{\rightarrow i}^{\,k}}\Biggl{)}\,$, (18)
where $\bar{y}_{\rightarrow i}^{\,k}\in\mathbb{R}^{n}$ is the desired message containing the $y$-features of all neighbors from node $i$. By the same argument as above, we can leave the constant term involving $x_{i}^{k}$ out in the optimization.
Note that this rule can be modified further if the edge weights $e_{ij}$ are not fixed over the whole algorithm and are not accessible for node $i$. In this case, one must add $(P_{ji},P_{ji}^{2})$ to the message.
MLP Integration.
For step size learning (both global and node-level), we can predict and use a new step size in each iteration. To achieve this, we insert each MLP that predicts a new step size into the first node update function of each iteration, before the $x$ variable is updated. In Algorithm 2.2, this is exactly between lines 5 and 6 of the pseudocode. The predicted step size is then used during an entire iteration.
Conversely, in the edge-level task, we predict the edge-weights $e_{ij}$ only once and keep them fixed during the whole $K=10$ unrolled iterations. The corresponding MLP is placed before the first iteration of Algorithm 2.2.
Algorithm 2.1 One iteration of decentralized, distributed ADMM
1:for $i\in\mathcal{V}$ in parallel do
2:
3:  Communicate $y_{i}^{k}$ and $\lambda_{i}^{k}$ to all neighbors $j\in\mathcal{N}(i)$
4:  Update local solution candidate $x$:
[…]
$+\hskip-2.84544pt\underset{j\in\mathcal{N}(i)\cup\{i\}}{\sum}\hskip-2.84544pt\left((\lambda_{j}^{k})^{T}(P_{ji}x_{i})+\frac{\alpha}{2}\|P_{ji}(x_{i}-x_{i}^{k})+y_{j}^{k}\|_{2}^{2}\right)\biggl{)}\,.$
5:  Communicate $x_{i}^{k+1}$ to all neighbors $j\in\mathcal{N}(i)$
6:  Update local deviation-variable $y$:
$y_{i}^{k+1}=\frac{1}{d_{i}+1}\sum_{j\in\mathcal{N}(i)\cup\{i\}}P_{ij}x_{j}^{k+1}\,.$
7:  Update local dual-variable $\lambda$:
$\lambda_{i}^{k+1}=\lambda_{i}^{k}+\alpha y_{i}^{k+1}\,.$
8:end for
Algorithm 2.2 One iteration of decentralized, distributed ADMM as 2-block GNN
1:for $(j,i)\in\mathcal{E}$ in parallel do
2:  $m_{ji}\leftarrow\left(P_{ji}\lambda_{j}^{k},P_{ji}y_{j}^{k}\right)$ $\triangleright$ message 1
3:end for
4:for $i\in\mathcal{V}$ in parallel do
5:  $\left(\bar{\lambda}_{\rightarrow i}^{k},\bar{y}_{\rightarrow i}^{k}\right)\leftarrow\sum_{(j,i)\in\mathcal{E}}m_{ji}$ $\triangleright$ aggregation 1
6:     ${v}_{i}\hskip 19.34766pt\leftarrow\left(x_{i}^{k+1},y_{i}^{k},\lambda_{i}^{k}\right),\,\text{where }\triangleright$ update 1
$x_{i}^{k+1}=\underset{x_{i}}{\text{arg min}}\,\,\biggl{(}f_{i}(x_{i})+\bigl{(}P_{ii}\lambda_{i}^{k}+$ $\bar{\lambda}_{\rightarrow i}^{\,k}+\,\alpha(P_{ii}y_{i}^{k}+\bar{y}_{\rightarrow i}^{\,k})\bigl{)}^{T}x_{i}$
$+\frac{\alpha}{2}\sum_{j\in\mathcal{N}(i)\cup\{i\}}\|P_{ji}(x_{i}-x_{i}^{k})\|_{2}^{2}\biggr{)}\,.$
7:end for
8:for $(j,i)\in\mathcal{E}$ in parallel do
9:  ${m}_{ji}\leftarrow P_{ij}x_{j}^{k+1}$ $\triangleright$ message 2
10:end for
11:for $i\in\mathcal{V}$ in parallel do
12:  $\bar{x}_{\rightarrow i}^{k+1}\leftarrow\sum_{(j,i)\in\mathcal{E}}m_{ji}$ $\triangleright$ aggregation 2
13:   ${v}_{i}\hskip 6.544pt\leftarrow\left(x_{i}^{k+1},y_{i}^{k+1},\lambda_{i}^{k+1}\right),\,\text{where }$ $\triangleright$ update 2
$y_{i}^{k+1}$ $=\frac{1}{d_{i}+1}\left(\bar{x}_{\rightarrow i}^{k+1}+P_{ii}x_{i}^{k+1}\right)\,$,
$\lambda_{i}^{k+1}$ $=\lambda_{i}^{k}+\alpha y_{i}^{k+1}\,.$
14:end for